OpenStudy (anonymous):

Solve the following equation for theta where 0 < or= theta < 2 pie sin ^2 theta = sin theta

OpenStudy (freckles):

\[\sin^2(\theta)-\sin(\theta)=0\] factor left hand side

OpenStudy (anonymous):

is that (sin theta) (sin theta)???

OpenStudy (freckles):

yes x^2 can be written as x*x

OpenStudy (anonymous):

so... now what?

OpenStudy (freckles):

factor

OpenStudy (freckles):

\[x^2-x \\ x(x-1)\] you try factoring \[\sin^2(\theta)-\sin(\theta)\] it is the same way as I did the expression of x

OpenStudy (freckles):

\[x^2=x \\ x^2-x=0 \\ x(x-1)=0 \\ \text{ set both factors equal to zero } \\ x=0 \text{ or } x-1=0 \\ x=0 \text{ or } x=1 \]

OpenStudy (freckles):

you try solving for sin(theta) now it is the same way

OpenStudy (freckles):

once you have solved for sin(theta) then you will easily be able to find theta

OpenStudy (anonymous):

hold on, don't leave yet

OpenStudy (anonymous):

sin =0 sin theta = 0

OpenStudy (anonymous):

I have no idea what to do next

OpenStudy (freckles):

well you have one equation correct but you are missing an equation

OpenStudy (freckles):

have you not factored \[\sin^2(\theta)-\sin(\theta)\] yet?

OpenStudy (freckles):

do you see both sin^2(theta) and sin(theta) have a common factor?

OpenStudy (freckles):

as you said sin^2(theta) can be written as sin(theta)*sin(theta) therefore sin(theta) and sin^2(theta) both have the common factor sin(theta)

OpenStudy (freckles):

so you can factor that out of the expression

OpenStudy (freckles):

\[\sin^2(\theta)-\sin(\theta)=0 \\ \sin(\theta)(\sin(\theta)-1)=0\] that is what you should get for the left hand side after factoring now set both factors equal to 0

OpenStudy (freckles):

you can use the unit circle to solve both equations

OpenStudy (anonymous):

wait hold on

OpenStudy (anonymous):

where did you get these 2 equations?

OpenStudy (freckles):

which two?

OpenStudy (anonymous):

wait nvm

OpenStudy (anonymous):

ok but how do you use the unit circle for this? i'm lost

OpenStudy (freckles):

do you know which equations you are solving?

OpenStudy (anonymous):

sin^2 theta - sin theta = 0

OpenStudy (freckles):

you need to factor the left hand side

OpenStudy (freckles):

I have actually done this for you

OpenStudy (anonymous):

I know x factor, but don't know how

OpenStudy (freckles):



OpenStudy (freckles):

now set both factors equal to 0

OpenStudy (freckles):

you have sin(theta)=0 or sin(theta)-1=0

OpenStudy (freckles):

find when sin(theta) is 0 on the unit circle find when sin(theta) is 1 on the unit circle

OpenStudy (freckles):

in other words you are looking for where the y-coordinate is 0 on the unit circle and also looking for when the y coordinate is 1 on the unit circle

OpenStudy (anonymous):

the answer is pie/2 and pie

OpenStudy (anonymous):

but how is it pie though???

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

got it

OpenStudy (freckles):

so you don't get that sin(pi)=0? also sin(0)=0 so I think you mean you have 0,pi/2,pi

OpenStudy (anonymous):

ok thx